OpenStudy (anonymous):

PUT INTO LOWEST TERMS 3/6 2/8 9/6 9/12 8/10 6/4 8/16a 2d/10

OpenStudy (anonymous):

1/2 1/4 3/2 3/4 yu see the pattern - just divide top and bottom by the common number

OpenStudy (anonymous):

what about 2d/10?

Directrix (directrix):

8/10 = 4/5 6/4 = 3/2 8/16a = 1/(2a) if the problem is 8/(16a) If the problem is (8/16) a, then (1/2) a = a/2 2d/10 = d/5

OpenStudy (anonymous):

and, cd/6d 6a/16b 3y/12xy 3r2/18r 10r/6s2 5yzr/10z

Directrix (directrix):

@pugbug --> post your answers for these and I will check them.

OpenStudy (anonymous):

c/6 3a/8b 1/4x 1r/6 5r/3s^2 yr/2

Directrix (directrix):

cd/6d = c/6 6a/16b = 3a/8b 3y/12xy = 1/4x 3r^2/18r = r/6 10r/6s^2 = 5r/3s^2 5yzr/10z = yr/2